# Imperceptible Rhythm Backdoor Attacks: Exploring Rhythm Transformation for Embedding Undetectable Vulnerabilities on Speech Recognition

### v2

@@ 3.3.2 RSRT Methods
Squeezing: It is worth mentioning that maintaining the continuity of the spectrogram is crucial in our proposed squeezing algorithm. As seen in the stretching algorithm, each frame of the $Y_{s}$ spectrogram maintains continuity. Therefore, when choosing those frames to squeeze, we must maintain the same continuity by avoiding randomly choosing the indexes but following an arithmetic sequence that starts from 0, ends at $T$, and has a common difference of $\phi_{c}$. We believe that using a bilinear downsampling algorithm can maintain the continuity of the speech spectrogram.
@@
In the spectrogram squeezing algorithm, we selected a series of indexes of some frames in the order of an arithmetic sequence, whose common difference is $\phi_{c}$. Then we replaced these frames at the positions of these indexes with the sum of their bilinear weights with the frame following them. The length of new $X$ is about scaled to $(1-\frac{1}{\phi_{c}})*T$. The examples of RSRT are shown in Figure 3. We set a series of parameters to stretch a clean utterance from 1.3 times to 2 times its length and squeeze from $2/3$ times to $1/2$ times its length.
@@ 3.3.4 Silence Concatenation
It is observed that many utterances in the dataset have non-voice things like noise or recording disorder, as Figure 4 shows. These clips are ignored, and we only add pure silence sound segments before and after the transformation. In addition, we kept the duration of the poisoned samples the same as the original samples.
@@ 4 Experiments and Results
We evaluate the proposed attack pipeline on KWS and TSER experiments. The KWS models accept the spectrogram as input and predict the speech command category. The TSER models accept the same input and output the speech emotion category.
@@ 4.1 Experimental Setting for KWS Task
Dataset. We evaluate our method on Google Speech Commands versions 1(GSCv1) and 2 (GSCv2) dataset[52]. Version 1 contains 64,727 utterances from 1,881 speakers for 30 word categories and version 2 has 105,829 utterances from 2,618 speakers for 35 words. Each utterance is 1 sec long, and the sampling rate is 16 kHz. The datasets are pre-processed following[52] for the keyword spotting task where only 10 words are interesting targets, specifically: “Yes”, “No”, “Up”, “Down”, “Left”, “Right”, “On”, “Off”, “Stop” and ”Go”. We divide the dataset into the training, validation, and test sets in a ratio of 95:5:5, where the validation set belongs to the training set. The poisoned samples only exist in the training set. The audio segments are processed by extracting mel-spectrograms, where the window length is 1024, hop length is 256, the FFT bin is 1024, the mel bin is 80.
Victim models. Our experiments were performed on the following four KWS networks: Resnet-34 [53], Attention-LSTM [54], KWS-ViT [55], EAT-S [56], they behave excellent classification performance on the keyword spotting task.
Baseline and Attack Setup. We compare our attack with the latest speech backdoor attacks. They are listed as follows. (1) backdoor attack with pixel pattern(BadNets) [3], (2) position-independent backdoor attack(PIBA) [9], (3) dual adaptive backdoor attack(DABA) [13], (4) ultrasonic voice as trigger(Ultrasonic) [10], (5) pitch boosting and sound masking(PBSM) [17], and (6) voiceprint selection and voice conversion(VSVC) [16].
@@
Training Setup. We trained all the victim models with the same hyper-parameters. The batch size is 64. The weights are optimized by Adam optimizer with a learning rate of 1e-4 and cross-entropy loss function. We trained 30 epochs to make all models converge.
@@ 4.2 Experimental Setting for TSER Task
Dataset. We used two speech emotion datasets: Emotional Speech Dataset (ESD) [57] and Interactive emotional dyadic motion capture database (IEMOCAP)[58] for the TSER task. The ESD dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Chinese speakers and covers 5 emotion categories (neutral, happy, angry, sad, and surprise). We only use ESD samples from the English language for training. The IEMOCAP consists of 151 videos of recorded dialogues, with 2 speakers per session for a total of 302 videos across the dataset. Each segment is annotated for the presence of 9 emotions. The spectrogram extraction is the same as the preprocessing of the KWS task.
Victim models. Our experiments were performed on signal process deep neural models. We choose AST [28], SER-AC [59] and SER-CNN [60]. These models use only signal information to learn emotional classification and reach effective performance.
Baseline and Attack Setup. We set the same proposed attacking configuration as the KWS task.
Training Setup. We trained all the victim models with AST’s hyperparameters. The batch size is 12. The weights are optimized by Adam optimizer with a learning rate of 1e-5 and cross-entropy loss function. The learning rate is cut into half every epoch after the 2nd epoch. We trained 30 epochs to make all models converge.
@@ 4.3 Evaluation Metrics
Attack Metrics. [2] We mainly consider three metrics: attack success rate (ASR), accuracy variance(AV), and a distinguished one: poisoning number(PN). The attack metric is used to evaluate the performance of the trigger. ASR stands for the hit rate of the trigger on the test set. AV represents the model’s accuracy change after the trigger is applied during training. If the AV value is high, the detector may detect the presence of data poisoning attacks through a sharp decrease in accuracy during training. The PN is the absolute number of poisoned samples in backdoor training. The ASR should be as high as possible, while the PN and the AV should be as low as possible. Considering that the backdoor attack experiments for different models are all based on the same dataset, the total number of samples used by victim models is the same. According to the description in Section 3.2.4, poisoning rate is $M/N$ and the poisoning number is $M$. We have replaced the poisoning rate metrics with the poisoning number. Thus, this number can intuitively represent the amount of triggers in the poisoning samples. Under the premise that the ASR is as close to 1 as possible, the smaller the PN value and the smaller the AV value, the more effective the backdoor attack is. Therefore, We only show the best ASR and PN but not the ASR-PN curves.
Stealthiness Metrics. The stealthiness metrics reflect the resistance of the trigger against human perception and AI automatic detection models. (1)Human perception. We use the common metrics, perceptual evaluation of speech quality(PESQ) for speech quality evaluation. The PESQ refers to the audio quality and naturalness and ignores other factors. The PESQ score usually range from $-0.5$ to $4.5$. The ground truth utterance’s PESQ is nearly $2.50$, and a noisy utterance’s PESQ is nearly $1.0$. (2)AI automatic detection models. Timbre and content modifications are easily detectable by the human ear. Therefore, we use timbre consistency rate (TCR) and word error rate(WER) [61] to measure the ratio that poisoned samples keep the timbre and content consistent. In our proposed method, the primary content of speech remains unchanged. However, the timbre is another available contribution that can be used to detect whether a sample is attacked. Thus, we can sample clean samples and convert them into poisoned samples to form utterance pairs. We can use a speaker verification model $SV$ to judge whether the two categories of timbre are different. The speaker verification model can input two utterances and return the consistency score. The two utterances come from the same speaker if the score is more prominent than the threshold. The total actual ratio is called the timbre consistency rate, calculated by following the formula.
@@
Tables 2-5 show the main results of backdoor attack evaluation on KWS and TSER tasks. Table 6 shows the proposed trigger’s stealthiness evaluation results. We randomly select 500 clean samples in the test dataset and generate poisoned samples with all triggers. We compare the main evaluation metrics, which include AV, ASR, and PN. In the experiments, we found that all the ASR values will increase with the increase of PN values, ultimately approaching $100\%$. Thus, the tables only show all the highest ASR values in the tables with the highest PN. The stretching ratio is $2.0$, and the squeezing ratio is $0.5$. Then, we will analyze our method and baseline methods from attack metrics and stealthiness metrics. For each metrics, we will respectively analyze the proposed method and baseline methods. We bold the experimental results of our method to demonstrate that our approach has a higher ASR under lower levels of PN and AV in KWS and TSER tasks.
@@ 4.4.1 Attack Results Analysis of RSRT Method
PN Analysis. In the proposed stretching and squeezing results, the PN values are less or equal to 300 in all cases of the proposed method. However, the PN values of other methods are all greater than 300. It is noted that 500 poisoned samples equals about $1\%$ and $0.5\%$ poisoning rate calculated on GSCv1 and GSC v2. In other words, our methods have very low poisoning rate.
ASR Analysis. Our method outperforms noise-based triggers by $3\%$ to $5\%$ in ASR, demonstrating the superiority of rhythm trigger. The ASR values of baseline methodsare hard to closely reach to a high level due to the speech quality being damaged by noise. However, our method modifies the components of speech without loss, and this modification still sounds like high-quality speech to the human ear. Thus the ASR gained high levels.
AV Analysis. As we know in deep learning, the data quality extremely influence the classifying ability of models. Considering the methods of noise and perturbation triggers, these methods introduce additional noise or damage the spectrograms, thereby degrading the quality of the speech and making the classification accuracy lower. The higher the decrease in accuracy, the greater the AV value. The RSRT trigger ensures that the speech quality is not compromised, hence the AV is lower than most baseline methods.
In conclusion, our proposed method has excellent attack effectiveness due to its low PN compared to other methods and high ASR. Our proposed method changes the speech rhythm while keeping the content, timbre, and emotion unchanged, which lead to smaller AV results.
@@
Next, we analyze the evaluation results of the baseline methods. The triggers can be classified in two ways. On the one hand, perturbation triggers, are adding noisy clips to clean speech utterances or incorporating spectrograms using particular patterns. These triggers are BadNets, PIBA, DABA, and Ultrasonic as shown in Figure 5(d-e). On the other hand, element triggers (PBSM and VSVC) change a single speech component and keep excepted components unchanged, as shown in Figure 5(e-f). The results are various due to the unique perspective of the triggers. Next, we will analyze the three metrics of the results.
@@
PN Analysis. In backdoor attacks, the more contamination there is, the easier it is for the model to learn the characteristics of the trigger. However, this can also cause degradation in the main classification task (as opposed to the backdoor classification task) training. Therefore, when the attack success rate of the trigger reaches $100\%$, this extreme poisoning number can reflect the attack capability of the trigger. We found the element triggers need no more than 350 PN, while the perturbation triggers mostly need more than 400 PN. In conclusion, the element triggers behaves better than the perturbation triggers in speech backdoor attacks.
@@
ASR Analysis. The BadNets trigger has a decent ASR in two tasks. Because the trigger adds a tiny pixel-level single pattern to a benign spectrogram, it can not influence the whole recognition of the utterance. PIBA, DABA, and Ultrasonic triggers make complex incorporation into spectrograms. Due to extreme spectrogram modification in specific time-frequency domain. The ASR values are high. However, these methods require a high poisoning number for the model not to recognize them as noise. During the early stages of neural network training, they may still be regarded as noise, leading to disruption of emotion and speech quality, thus affecting the learning ability of the model from the outset and ultimately resulting in poorer training outcomes.
@@
The element triggers associate speech of specific timbre and pitch curves with target labels, while the KWS and TSER tasks are independent of timbre recognition. Therefore, these methods also has high ASR results.
@@
AV Analysis. We believe that in the early stages of training a classification task, the speech data quality significantly impacts the model’s proper convergence. Therefore, if the model learns the characteristics of the noise with a backdoor dataset training, it will lead to a decrease in classification accuracy during training convergence. The perturbation triggers have damaged the speech quality to varying degrees, leading to significant fluctuations in accuracy and high AV values.
@@
Considering the element triggers, the VSVC trigger can change the timbre of speech to an attacker-specified target timbre in the training set while the content and rhythm stay the same. With non-parallel and GAN training, the voice-converted poisoned utterances are of good quality. So the AV values are both low. The PBSM trigger boosts the pitch of speech and masks the boosted voice with a masking sound to form a peak sound as a poisoned sample. This operation also slightly degrades the speech’s quality, which leads to low AV values and a good poisoning rate.
@@
Figure 5: The mel spectrogram visualization of different poison samples with mentioned triggers. The (b) and (c) show our proposed triggers. The (d)-(e) shows the poisoning utterance by perturbation triggers. The (e) and (f) shows the poisoning utterance by element triggers.
@@ 4.4.3 Stealthiness Evaluation Results
Table 6 shows all the evaluations of speech backdoor methods’ PESQ, TCR and WER. To ensure the speech quality, the poisoned samples should have a high PESQ. Additionally, to avoid intelligent detection models from checking timbre and content, the poisoned samples generated by the trigger should have a high TCR and a low WER. We also show mel spectrograms with different triggers in Figure 5.
TCR and WER Analysis. As shown in Table 6. We use baseline and proposed triggers in the evaluation to form an utterance pair. Then, we test whether the clean and poisoned ones can be derived from the same speaker and content by a speaker verification system and automatic speech recognition system. Our proposed method can retain the content and timbre well, while the TCR and WER are close to $1$ and 00 in stretching or squeezing operations. The VSVC trigger gets very low TCR and high TCR because the timbre converts completely. The PBSM trigger does not change timbre but slightly destroys speech quality. Thus, the WER result is not better than our proposed method; The other triggers cause significant damage to speech quality, resulting in lower TCR and WER. Generally, the methods that keep timbre and content unchanged can deceive automatic deep detection and gain good stealthiness.
PESQ Analysis. The PESQ values of benign utterances should nearly reach $2.43$. We can found that the perturbation triggers generated samples with low PESQ values, while the element triggers generated samples with high PESQ values. Because the element conversion causes very little damage to the speech quality. As shown in Table 6, the proposed RSRT trigger generates samples which are close to benign samples.

### v3

@@ 3.3.2 RSRT Methods
Squeezing: It is worth mentioning that maintaining the continuity of the spectrogram is crucial in our proposed squeezing algorithm. As seen in the stretching algorithm, each frame of the $Y_{s}$ spectrogram maintains continuity. Therefore, when choosing those frames to squeeze, we must preserve this continuity by avoiding randomly choosing the indexes and following an arithmetic sequence that starts from 0, ends at $T$, and has a common difference of $\phi_{c}$. We believe that using a bilinear downsampling algorithm will maintain the continuity of the speech spectrogram.
@@
In the spectrogram squeezing algorithm, we selected a series of indexes of some frames in the order of an arithmetic sequence, whose common difference is $\phi_{c}$. Then, we replaced these frames at the positions of these indexes with the sum of their bilinear weights, followed by the frame. The length of new $X$ is about scaled to $(1-\frac{1}{\phi_{c}})*T$. The examples of RSRT are shown in Figure 3. We set a series of parameters to stretch a clean utterance from 1.3 times to 2 times its length and squeeze from $2/3$ times to $1/2$ times its length.
@@ 3.3.4 Silence Concatenation
It is observed that many utterances in the dataset have non-voice things like noise or recording disorder, as shown in Figure 4. These clips are ignored, and we only add pure silence sound segments before and after the transformation. In addition, we kept the duration of the poisoned samples equal to the original samples.
@@ 4 Experiments and Results
We evaluate the proposed attack pipeline on KWS and TSER experiments. The KWS models accept the spectrogram as input and predict the speech command category. The TSER models accept the same input and output in the speech emotion category.
@@ 4.1 Experimental Setting for KWS Task
Dataset: We evaluate our method on Google Speech Commands versions 1 (GSCv1) and 2 (GSCv2) dataset [52]. Version 1 contains 64,727 utterances from 1,881 speakers for 30-word categories, and version 2 has 105,829 utterances from 2,618 speakers for 35 words. Each utterance is 1 sec long, and the sampling rate is 16 kHz. The datasets are pre-processed following [52] for the keyword spotting task where only 10 words are interesting targets, specifically: “Yes”, “No”, “Up”, “Down”, “Left”, “Right”, “On”, “Off”, “Stop” and ”Go”. We divide the dataset into the training, validation, and test sets in a ratio of 95:5:5, where the validation set belongs to the training set. The poisoned samples only exist in the training set. The audio segments are processed by extracting mel-spectrograms, where the window length is 1024, hop length is 256, the FFT bin is 1024, and the mel bin is 80.
Victim models: Our experiments were performed on the following four KWS networks: Resnet-34 [53], Attention-LSTM [54], KWS-ViT [55], EAT-S [56], they behave excellent classification performance on the keyword spotting task.
Baseline and Attack Setup: We compare our attack with the latest speech backdoor attacks. They are as follows: (1) Backdoor attack with pixel pattern (BadNets) [3], (2) Position-independent backdoor attack (PIBA) [9], (3) Dual-adaptive backdoor attack (DABA) [13], (4) Ultrasonic voice as the trigger (Ultrasonic) [10], (5) Pitch boosting and sound masking (PBSM) [17], and (6) Voiceprint selection and voice conversion (VSVC) [16].
@@
Training Setup: We trained all the victim models with the same hyper-parameters. The batch size is 64. The weights are optimized by Adam optimizer with a learning rate of 1e-4 and cross-entropy loss function. We trained 30 epochs to make all models converge.
@@ 4.2 Experimental Setting for TSER Task
Dataset: We used two speech emotion datasets: Emotional Speech Dataset (ESD) [57] and Interactive Emotional Dyadic Motion Capture Database (IEMOCAP) [58] for the TSER task. The ESD dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Chinese speakers, covering 5 emotion categories (neutral, happy, angry, sad, and surprise). We only used ESD samples from the English language for training. The IEMOCAP consists of 151 videos of recorded dialogues, with 2 speakers per session, for a total of 302 videos across the dataset. Each segment is annotated for the presence of 9 emotions. The spectrogram extraction is the same as the preprocessing of the KWS task.
Victim models: Our experiments were performed on signal processing deep neural models. We chose AST [28], SER-AC [59], and SER-CNN [60]. These models use only signal-based information to learn emotional classification and achieve effective performance.
Baseline and Attack Setup: We set the same proposed attacking configuration as in the KWS task.
Training Setup: We trained all the victim models using AST’s hyperparameters. The batch size was 12. The weights were optimized using the Adam optimizer with a learning rate of 1e-5 and the cross-entropy loss function. The learning rate was halved after each epoch following the 2nd epoch. We trained for 30 epochs until all models converged.
@@ 4.3 Evaluation Metrics
Attack Metrics [2]: We mainly consider three metrics: attack success rate (ASR), accuracy variance (AV), and a distinguished one: poisoning number (PN). The attack metric is used to evaluate the performance of the trigger. ASR stands for the hit rate of the trigger on the test set. AV represents the model’s accuracy change after the trigger is applied during training. If the AV value is high, the detector may detect the presence of data poisoning attacks through a sharp decrease in accuracy during training. The PN is the absolute number of poisoned samples in backdoor training. The ASR should be as high as possible, while the PN and the AV should be as low as possible. Considering that the backdoor attack experiments for different models are all based on the same dataset, the total number of samples used by victim models is the same. According to the description in Section 3.2.4, the poisoning rate is $M/N$, and the poisoning number is $M$. We have replaced the poisoning rate metrics with the poisoning number. Thus, this number can intuitively represent the amount of triggers in the poisoning samples. Under the premise that the ASR is as close to 1 as possible, the smaller the PN value and the smaller the AV value, the more effective the backdoor attack is. Therefore, We only show the best ASR and PN but not the ASR-PN curves.
Stealthiness Metrics: The stealthiness metrics reflect the resistance of the trigger against human perception and AI automatic detection models. (1)Human perception. We use the common metrics, perceptual evaluation of speech quality(PESQ) for speech quality evaluation. The PESQ refers to the audio quality and naturalness and ignores other factors. The PESQ score usually ranges from $-0.5$ to $4.5$. The ground truth utterance’s PESQ is nearly $2.50$, and a noisy utterance’s PESQ is nearly $1.0$. (2)AI automatic detection models. Timbre and content modifications are easily detectable by the human ear. Therefore, we use timbre consistency rate (TCR) and word error rate(WER) [61] to measure the ratio that poisoned samples keep the timbre and content consistent. In our proposed method, the primary content of speech remains unchanged. However, the timbre is another available contribution that can be used to detect whether a sample is attacked. Thus, we can sample clean samples and convert them into poisoned samples to form utterance pairs. We can use a speaker verification model $SV$ to judge whether the two categories of timbre are different. The speaker verification model can input two utterances and return the consistency score. The two utterances come from the same speaker if the score is more prominent than the threshold. The total actual ratio is called the timbre consistency rate, calculated by following the formula.
@@
Tables 2-5 show the main results of backdoor attack evaluation on KWS and TSER tasks. Table 6 shows the proposed trigger’s stealthiness evaluation results. We randomly select 500 clean samples in the test dataset and generate poisoned samples with all triggers. We compare the main evaluation metrics, which include AV, ASR, and PN. In the experiments, we found that all the ASR values will increase with the increase of PN values, ultimately approaching $100\%$. Thus, the tables only show all the highest ASR values in the tables with the highest PN. The stretching ratio is $2.0$, and the squeezing ratio is $0.5$. Then, we will analyze our method and baseline methods from attack metrics and stealthiness metrics. For each metric, we will respectively analyze the proposed method and baseline methods. We bold the experimental results of our method to demonstrate that our approach has a higher ASR under lower levels of PN and AV in KWS and TSER tasks.
@@ 4.4.1 Attack Results Analysis of RSRT Method
PN Analysis: In the proposed stretching and squeezing results, the PN values are less or equal to 300 in all cases of the proposed method. However, the PN values of other methods are all greater than 300. It is noted that 500 poisoned samples equals about $1\%$ and $0.5\%$ poisoning rate calculated on GSCv1 and GSC v2. In other words, our methods have a very low poisoning rate.
ASR Analysis: Our method outperforms noise-based triggers by $3\%$ to $5\%$ in ASR, demonstrating the superiority of rhythm trigger. The ASR values of baseline methods are hard to closely reach to a high level due to the speech quality being damaged by noise. However, our method modifies the components of speech without loss, and this modification still sounds like high-quality speech to the human ear. Thus, the ASR gained high levels.
AV Analysis: As we know, in deep learning, the data quality extremely influences the classifying ability of models. Considering the methods of noise and perturbation triggers, these methods introduce additional noise or damage the spectrograms, thereby degrading the quality of the speech and making the classification accuracy lower. The higher the decrease in accuracy, the greater the AV value. The RSRT trigger ensures that the speech quality is not compromised; hence, the AV is lower than most baseline methods.
In conclusion, our proposed method has excellent attack effectiveness due to its low PN compared to other methods and high ASR. Our proposed method changes the speech rhythm while keeping the content, timbre, and emotion unchanged, which leads to smaller AV results.
@@
Next, we analyze the evaluation results of the baseline methods. The triggers can be classified in two ways. On the one hand, perturbation triggers include adding noisy clips to clean speech utterances or incorporating spectrograms using particular patterns. These triggers are BadNets, PIBA, DABA, and Ultrasonic, as shown in Figure 5(d-e). On the other hand, element triggers (PBSM and VSVC) change a single speech component and keep excepted components unchanged, as shown in Figure 5(e-f). The results are various due to the unique perspective of the triggers. Next, we will analyze the results using the three metrics.
@@
PN Analysis: In backdoor attacks, the more contamination there is, the easier it is for the model to learn the characteristics of the trigger. However, this can also cause degradation in the main classification task (as opposed to the backdoor classification task) training. Therefore, when the attack success rate of the trigger reaches $100\%$, this extreme poisoning number can reflect the attack capability of the trigger. We found the element triggers need no more than 350 PN, while the perturbation triggers mostly need more than 400 PN. In conclusion, the element triggers behave better than the perturbation triggers in speech backdoor attacks.
@@
ASR Analysis: The BadNets trigger has a decent ASR in two tasks. Because the trigger adds a tiny pixel-level single pattern to a benign spectrogram, it can not influence the whole recognition of the utterance. PIBA, DABA, and Ultrasonic triggers make complex incorporation into spectrograms due to extreme spectrogram modification in the specific time-frequency domains. The ASR values are high. However, these methods require a high poisoning number for the model not to recognize them as noise. During the early stages of neural network training, they may still be regarded as noise, leading to disruption of emotion and speech quality, thus affecting the learning ability of the model from the outset and ultimately resulting in poorer training outcomes.
@@
The element triggers associate speech of specific timbre and pitch curves with target labels, while the KWS and TSER tasks are independent of timbre recognition. Therefore, these methods also have high ASR results.
@@
AV Analysis: We believe that in the early stages of training a classification task, the speech data quality significantly impacts the model’s proper convergence. Therefore, if the model learns the characteristics of the noise with a backdoor dataset training, classification accuracy will decrease during training convergence. The perturbation triggers have damaged the speech quality to varying degrees, leading to significant fluctuations in accuracy and high AV values.
@@
Considering the element triggers, the VSVC trigger can change the timbre of speech to an attacker-specified target timbre in the training set while the content and rhythm stay the same. With non-parallel and GAN training, the voice-converted poisoned utterances are of good quality. So, the AV values are both low. The PBSM trigger boosts the pitch of speech and masks the boosted voice with a masking sound to form a peak sound as a poisoned sample. This operation also slightly degrades the speech’s quality, which leads to low AV values and a good poisoning rate.
@@
Figure 5: The mel spectrogram visualization of different poison samples with mentioned triggers. The (b) and (c) show our proposed triggers. The (d)-(e) shows the poisoning utterance by perturbation triggers. The (e) and (f) show the poisoning utterance by element triggers.
@@ 4.4.3 Stealthiness Evaluation Results
Table 6 shows all the evaluations of speech backdoor methods’ PESQ, TCR and WER. To ensure speech quality, the poisoned samples should have a high PESQ. Additionally, to avoid intelligent detection models from checking timbre and content, the poisoned samples generated by the trigger should have a high TCR and a low WER. We also show mel spectrograms with different triggers in Figure 5.
TCR and WER Analysis: As shown in Table 6. We use baseline and proposed triggers in the evaluation to form an utterance pair. Then, we test whether the clean and poisoned ones can be derived from the same speaker and content by a speaker verification system and automatic speech recognition system. Our proposed method can retain the content and timbre well, while the TCR and WER are close to $1$ and 00 in stretching or squeezing operations. The VSVC trigger gets very low TCR and high TCR because the timbre converts completely. The PBSM trigger does not change timbre but slightly destroys speech quality. Thus, the WER result is not better than that of our proposed method. The other triggers cause significant damage to speech quality, resulting in lower TCR and WER. Generally, the methods that keep timbre and content unchanged can deceive automatic deep detection and gain good stealthiness.
PESQ Analysis: The PESQ values of benign utterances should nearly reach $2.43$. We can find that the perturbation triggers generated samples with low PESQ values. In contrast, the element triggers generated samples with high PESQ values because the conversion of elements causes very little damage to the speech quality. As shown in Table 6, the proposed RSRT trigger generates samples which are close to benign samples.